Check my workCheck My Work button is now enabled1

Item 2

Item 2 10 points Item Skipped

The number of blood tests performed and the related costs over the last nine months in Brentline Hospital are given below:

  Month Blood Tests
Performed
Blood
Test Costs
  January 4,625     $ 81,516
  February 7,750     $ 119,450
  March 4,125     $ 74,504
  April 3,680     $ 68,748
  May 3,915     $ 72,178
  June 3,000     $ 60,550
  July 3,455     $ 65,688
  August 4,725     $ 81,704
  September 4,785     $ 83,164

Required:

1. Using the high–low method, estimate the cost formula for blood tests. (Round cost per unit answer to 2 decimal places.)

2. Using the cost formula you derived above, what blood test costs would you expect to be incurred during a month in which 4,270 blood tests are performed? (Round your intermediate calculations answer to 2 decimal places.)
Answer #1

Requirement 1

Y= $          23,350.00 + $                12.40 X

Fixed cost =23500 and variable cost = 12.40 per unit

Requirement 2

Total cost for 4270 blood tests=$81,878.00

Working

Cost No. of activities
A High Level $        119,450.00 7750
B Low Level $          60,550.00 3000
C=A-B Difference $          58,900.00 $              4,750.00
A Cost difference $          58,900.00
B No. of activities difference 4750
C=A/B Variable cost per unit $ 12.40
High Level Low Level
A No. of activities 7750 3000
B Variable cost per unit $              12.4000 $               12.4000
C=AxB Total Variable cost $          96,100.00 $            37,200.00
D Total cost $        119,450.00 $            60,550.00
E=D-C Total fixed cost $          23,350.00 $            23,350.00